OpenStudy (anonymous):

More statistics. Binomial Probability \[P(r)=\frac{n!}{r!(n-r)!}p^rq^{n-r}=C_{n,r}p^rq^{n-r}\] The problem states 10% of adults deliberately do a one time fling (purchase clothing, wear 'em to an event, and return 'em). In a group of 7 adults what is the probability that anyone has done a one time fling? I figured that: n=7 r=0 I plugged those into the equation: \[P(0)=\frac{7!}{0!(7-0)!}p^0q^{7-0}=C_{7,0}p^0q^{7-0}\] What about p and q? The probability of success (p) is 0.1 (1% of people do one time flings)? and q would be 0.9? Would that be right?

OpenStudy (kropot72):

\[P(0\ out\ of\ 7)=7C0\times 0.1^{0}\times 0.9^{7}=0.9^{7}\] \[P(at\ least\ 1\ out\ of\ 7)=1.0000-0.9^{7}=\ ?\]

OpenStudy (kropot72):

\[q=(1-p)\]